Suppose we observe the following rates: 1R1 = 9%, 1R2 = 11%. If the unbiased expectations theory of the term structure of interest rates holds, what is the 1-year interest rate expected one year from now, E(2r1)? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

1 year interest rate expected one year from now be r

year 2 interest rate is a product of year one interest now and 1 year interest rate expected one year from now

(1+r) (1+0.09) = (1+0.11)2

1+r = 1.1304

r = 0.1304 or 13.04%